# Advancements in Visual Language Models for Remote Sensing: Datasets, Capabilities, and Enhancement Techniques

### v2

@@ 5.2 Advancements in Conversational Vision-Language Models
Leveraging the existing API of pre-trained LLMs, conversational VLMs extract features from image data using a visual encoder. These features are then transformed into a format that LLMs can interpret, enabling the completion of various visual language tasks. To enhance task performance, researchers meticulously designed each component, which is described in detail below. Figure 8 clearly shows the model details of conversational VLMs.
@@
Image Encoder Table 3 presents the visual encoders commonly used in current VLM implementations. Most studies utilizes the standard CLIP image encoder [17], but RSGPT [30] and SkyEyeGPT [44] have explored other variants, EVA-CLIP encoder [88], which is trained with improved training techniques.
Additionally, EarthGPT[45] adopt the vision transformer (ViT) as the image encoder to extract multi-layer features, which is beneficial to capture subtle differences in RS images and provides a more comprehensive understanding of images. Then the CNN backbone is designed as the image encoder to integrate multi-scale local details into visual representation. GeoChat [25], in addition to using CLIP-ViT(L-14) as the visual backbone network, interpolates the positional encoding in the transformer-based CLIP model with high-resolution 512x512 to understand details presented in remote sensing imagery (e.g., small objects and object details).
@@
Table 3: Architecture of conversational VLMs, $\ast$: frozen, $\blacktriangle$: fine-tuning, SC=Scene Classification, VQA=Visual Question Answer, VG=Visual Grounding, IC=Image Captioning, OD=Object Detection, IR=Image Retrieval, RSICC=Remote Sensing Image Change Captioning. For each model, we indicate the LLM used in its best configuration as shown in the original paper.
Model Image Encoder Text Encoder Capabilitie
RS-LLaVA [89] ViT-L(CLIP) Vicuna-v1.5-13B IC,VQA
H2RSVLM [46] ViT-L (CLIP) Vicuna-v1.5-7B SC,VQA,VG
SkySenseGPT [58] ViT-L (CLIP) Vicuna-v1.5-7B SC,IC,VQA,OD
GeoChat [25] ViT-L (CLIP) Vicuna-v1.5-7B SC,VQA,VGRegion Caption
RSGPT [30] ViT-G (EVA) Vicuna-v1.5-7B IC,VQA
SkyEyeGPT [44] ViT-G(EVA) LLaMA2-7B IC,VQA,VG
RS-CapRet [90] ViT-L(CLIP) LLaMA2-7B IC,IR
LHRS-Bot [54] ViT-L (CLIP) LLaMA2-7B SC,VQA,VG
EarthGPT [45] ViT-L (DINOv2) ConvNeXt-L (CLIP) LLaMA2-7B SC,VQA,IC,VG,OD
Liu et al. [81] ViT-B(CLIP) GPT-2 RSICC
@@
Text Encoder Training large language models (LLMs) from scratch is highly costly. Therefore, conversational visual language models (VLMs) utilize pre-trained LLMs as text encoders. These LLMs process visual elements extracted by the visual encoder and converted by middleware, along with the researcher’s questions or language instructions, to generate the desired results. Currently, the academic community is focusing on the LLaMA [21] and Vicuna [69] series, which serve as text encoders for most conversational LLMs. Table 3 provides a summary of these models. Common approaches include directly using LLMs or efficiently fine-tuning them with Low-Rank Adaptation (LoRA) [91]. Liu et al. [81] designed a complex image classifier and proposed a multi-prompt learning strategy to generate multiple learnable prompt embeddings as prefixes for LLMs to enhance their utilization for the RSICC task. RS-LLaVA [89] and RSGPT [30] demonstrated the capabilities of the proposed architecture using two different LLMs (vicuna-7B and vicuna-13B). The results indicated that fine-tuning larger language models leads to higher performance but at the cost of longer training times and higher computational resource requirements.
@@
Vision-Language Alighment In the architecture of conversational Vision-Language Models (VLMs), the middle connection layer plays a crucial role. It acts as a bridge between remote sensing images and Large Language Models (LLMs), projecting image information into a space that LLMs can comprehend, enabling them to understand complex remote sensing (RS) tasks. A common approach to bridge the modal gap is to use Multi-Layer Perceptrons (MLPs), which employ single linear layer [44, 45] or double layers [46, 58, 89] to project visual tokens and align feature dimensions with word embeddings. Additionally, some works focus on designing connectors with intricate structures, offering valuable insights for future research. SkyEyeGPT [44], to balance the important visual information processing of high-resolution remote sensing images by LLMs and the high resource requirements, uses the method by Chen et al. [92] to directly concatenate 4 adjacent visual tokens, reducing the number of patches by 4 times.
@@
One reason current general VLMs perform poorly in the remote sensing field is that the visual encoder of LLMs usually uses CLIP, but CLIP’s ViT lacks spatial awareness, affecting VLM’s spatial understanding ability [93]. To address this issue, H2RSVLM [46] and SkySenseGPT [58] enhance the visual encoder’s perceptual ability at the source by creating higher quality, large-scale, and detailed image-text pairs in remote sensing datasets. EarthGPT [45] expands the input images to include SAR and infrared images. Furthermore, some research endeavors opt for designing elaborate visual language connection structures to enhance the multimodal perception capabilities of VLM. GeoChat [25] constructs bounding boxes for text representations to express their spatial coordinates, enabling it to perform object grounding or accept region inputs, thus having reasoning capabilities for remote sensing images. RSGPT [30], under the guidance of InstructBLIP [94], inserts an annotation-aware query transformer (Q-Former) to enhance the alignment of visual and textual features. Additionally, the annotation-aware mechanism interacts with query embeddings through self-attention layers, allowing RSGPT to extract comprehensive visual features crucial for RS visual language tasks. LHRS-Bot [54] introduces a set of learnable queries for each level of image features. These queries aim to summarize the semantic information of each level of visual features through a series of stacked cross-attention and MLP layers.
@@
Model Enhancement Technique
@@
Txt2Img [95] Employ modern Hopfield layers for hierarchical prototype learning of text and image embeddings.
@@
CPSeg [96] Utilize "chain of thought" process that leverages textual information associated with the image.
@@
SHRNet [97] Utilize a hash-based architecture to extract multi-scale image features.
@@
MGeo [98] Introduce the geographic environment (GC) extracting multi-modal correlations for accurate query-POI matching.
@@
GeoCLIP [99] Via an image-to-GPS retrieval approach by explicitly aligning image features with corresponding GPS locations.
@@
SpectralGPT [100] Provide a foundational model for the spectral data based on the MAE[101] architecture.
@@
TEMO [102] Capture long-term dependencies from description sentences of varying lengths, generating text modal features for each category.
5.3 Others
In addition to the aforementioned approaches, as shown in Table 4, several other advancements have been made to address domain-specific tasks in remote sensing. Txt2Img [95] employs modern Hopfield layers for hierarchical prototype learning of text and image embeddings, achieving notable performance in generating text descriptions for remote sensing images. CPSeg [96] enhances image segmentation by integrating a novel "chain of thought" process that leverages textual information associated with the image. SHRNet [97] utilizes a hash-based architecture to extract multi-scale image features, while QRN improves alignment efficiency and enhances the visual-spatial reasoning ability of remote sensing visual question answering (VQA) systems. MGeo [98], a multi-modal geographic language model, introduces the geographic environment (GC) as a new modality, effectively extracting multi-modal correlations for accurate query-POI matching. GeoCLIP [99], with its independently designed Location Encoder, is the first to tackle worldwide geo-localization via an image-to-GPS retrieval approach by explicitly aligning image features with corresponding GPS locations. SpectralGPT [100], based on the MAE[101] architecture, provides a foundational model for the spectral data domain, supporting various downstream tasks. TEMO [102] enhances the word embedding component of VLMs, capturing long-term dependencies from description sentences of varying lengths, generating text modal features for each category, and employing a masking strategy to prevent overfitting. This method is particularly effective for few-shot object detection (FSOD) in text-modal knowledge-oriented remote sensing imagery.
@@ 5.4 Performance Comparison
In this section, we present a comparative analysis of various Visual-Language Model (VLM) pre-training methods, focusing on their performance across common datasets. It is important to note that we have prioritized performance evaluations on widely recognized datasets, which has resulted in the exclusion of methods that are tailored for specific datasets.
Table  5 summarizes the evaluation results for five prevalent image classification tasks, highlighting the best-performing VLMs. Due to the differing implementations of VLM pre-training methods, it is crucial to identify the standout models. Among those utilizing a contrastive approach, RemoteCLIP [24] emerges as the top performer, while the conversational approach is epitomized by SkySenseGPT [58], which consistently demonstrates superior performance. Overall, our findings suggest that models employing conversational techniques tend to outperform their contrastive counterparts.
Table 5: Performance of VLM pre-training methods on common remote sensing datasets, the best comprehensive performance was deepened separately. $\ast$: few-shot prediction.
@@
Table 6: Performance of VLM pre-training methods on visual question answer task, the best comprehensive performance was deepened separately. $\ast$: few-shot prediction.
@@
SHRNet [97] 85.85 85.39 84.46
MQVQA [38] - - 82.18
Table 7: Performance of VLM pre-training methods on image caption task, the best comprehensive performance was deepened separately.
@@
Syndney- Captions [75] CRSR [82] 79.94 74.4 69.87 66.02 41.50 74.88 289.00
@@
In Tables  6 and  7, we delve deeper into two frequently encountered visual-language tasks: Visual Question Answering (VQA) and Image Classification (IC). The results clearly illustrate that the conversational model, SkySenseGPT [58], maintains its exceptional performance across both tasks. Notably, none of the contrastive VLMs proved effective for the VQA task, further underscoring the advantages of conversational methods. Additionally, in the IC task, the prevalence of conversational VLMs exceeds that of their contrastive counterparts.
In conclusion, our analysis indicates that conversational VLMs not only demonstrate greater versatility across a range of tasks but also deliver superior performance overall. Consequently, we advocate for the adoption of conversational model architectures as the preferred choice for future developments in visual-language modeling.
@@ 6 Conclusions and Future Work
This review provides a systematic overview of Vision-Language Models (VLM) in the field of remote sensing. It introduces milestone works in the development of VLM itself, discusses the current research status of VLM in remote sensing, outlines the tasks involved, and offers a brief comparison of different VLM methods in various remote sensing applications. Overall, VLM has once again facilitated advancements in data processing technologies within the field of remote sensing. Current work has clearly demonstrated the potential of VLM applications in this area. However, there are still some shortcomings in current VLM technologies and their applications in remote sensing. Future work could consider the following directions:
  • Addressing regression problems. Currently, the mainstream approach in VLM involves using a tokenizer to vectorize text, then aligning visual information with the text word vectors before feeding them into the LLM model. This approach results in the model being insensitive to numerical values. For example, the number 100 may be tokenized as ’1’, ’00’, or it could be split into ’10’, ’0’, or ’100’. This leads to suboptimal performance when the current framework is applied to certain regression tasks. Future research on how to integrate regression problems into the current framework is a topic worth exploring.
  • Aligning with the structural characteristics of remote sensing images. Current VLM efforts in the remote sensing field largely follow the overarching framework established in the computer vision domain, without fully leveraging the unique characteristics of remote sensing images. For instance, remote sensing images come in various types, such as SAR images, HSI, and radar images, each with distinct features. Designing feature extractors that cater to these data characteristics is essential for maximizing the advantages of the images. However, in the current remote sensing VLM work, visual encoders primarily rely on pre-trained feature extractors based on RGB images, which do not fully align with remote sensing images. Additionally, there is a lack of research considering the fusion of different modalities of remote sensing images to enhance the performance of VLM.
  • Multimodal output. Existing VLM models in the remote sensing field typically output only text. While they can handle traditional tasks like image classification, they struggle with dense prediction tasks such as segmentation, where text output is insufficient. Therefore, researching multimodal outputs—allowing VLM models to also output images or even videos—is a highly valuable direction. This could facilitate the integration of various application needs within the remote sensing domain into a unified framework.

### v3

@@ 5.2 Advancements in Conversational VLMs
Leveraging the existing API of pre-trained LLMs, conversational VLMs extract features from image data using a visual encoder. These features are then transformed into a format that LLMs can interpret, enabling the completion of various vision–language tasks. To enhance task performance, researchers meticulously design each component, which is described in detail below. Figure 10 clearly shows the model details of conversational VLMs.
@@
Image Encoder: Table 3 presents the visual encoders commonly used in current VLM implementations. Most studies utilize the standard CLIP image encoder [22], but RSGPT [35] and SkyEyeGPT [54] have explored other variants, such as EVA-CLIP encoder [98], which is trained with improved training techniques.
Additionally, RS-LLaVa [99] adapts the LLaVA framework for the remote sensing domain by using a domain-specific image encoder and training it on specialized RS datasets. Through instruction tuning and precise visual–language alignment, RS-LLaVa gains a deeper understanding of geospatial semantics, allowing it to provide contextually relevant, detail-oriented responses tailored to the complexities of remote sensing imagery. EarthGPT [55] employs a Vision Transformer as its image encoder to extract multi-layered features that capture subtle differences in RS images. It then integrates a CNN backbone to incorporate multi-scale local details into the visual representation, delivering more comprehensive analytical capabilities. GeoChat [30] builds on the CLIP-ViT(L-14) backbone, interpolating positional encodings at higher resolutions (512 × 512) to better discern fine-grained details. This approach enhances its capacity to identify small objects and subtle characteristics in remote sensing imagery.
@@
Table 3: Architecture of conversational VLMs. $\ast$: frozen; $\blacktriangle$: fine-tuning; SC = Scene Classification; VQA = Visual Question Answering; VG = Visual Grounding; IC = Image Captioning; OD = object detection; IR = Image Retrieval; RSICC = Remote Sensing Image Change Captioning. For each model, we indicate the LLM used in its best configuration as shown in the original paper.
Model Year Image Encoder Text Encoder Capability
RS-LLaVA [99] 2024 ViT-L(CLIP)  Vicuna-v1.5-13B  IC, VQA
H2RSVLM [56] 2024 ViT-L (CLIP)  Vicuna-v1.5-7B  SC, VQA, VG
SkySenseGPT [68] 2024 ViT-L (CLIP)  Vicuna-v1.5-7B  SC, IC, VQA, OD
GeoChat [30] 2024 ViT-L (CLIP)  Vicuna-v1.5-7B  SC, VQA, VG Region Caption
RSGPT [35] 2023 ViT-G (EVA)  Vicuna-v1.5-7B  IC, VQA
SkyEyeGPT [54] 2024 ViT-G (EVA)  LLaMA2-7B  IC, VQA, VG
RS-CapRet [100] 2024 ViT-L (CLIP)  LLaMA2-7B  IC, IR
LHRS-Bot [64] 2024 ViT-L (CLIP)  LLaMA2-7B  SC, VQA, VG
EarthGPT [55] 2024 ViT-L (DINOv2)  ConvNeXt-L (CLIP)  LLaMA2-7B  SC, VQA, IC, VG, OD
Liu et al. [91] 2023 ViT-B(CLIP)  GPT-2  RSICC
@@
Text Encoder: Training large language models (LLMs) from scratch is highly costly. Therefore, conversational Vision–Language Models (VLMs) utilize pre-trained LLMs as text encoders. These LLMs process visual elements extracted by the visual encoder and converted by middleware, along with the researcher’s questions or language instructions, to generate the desired results. Currently, the academic community is focusing on the LLaMA [26] and Vicuna [78] series, which serve as text encoders for most conversational LLMs. Table 3 provides a summary of these models. Common approaches include directly using LLMs or efficiently fine-tuning them with Low-Rank Adaptation (LoRA) [101]. Liu et al. [91] designed a complex image classifier and proposed a multi-prompt learning strategy to generate multiple learnable prompt embeddings as prefixes for LLMs to enhance their utilization for the RSICC task. In addition, RS-LLaVA [99] applies the LLaVA framework to remote sensing data, incorporating a Vicuna-v1.5-13B LLM along with domain-specific image encoders and specialized training corpora, enabling the model to better capture geospatial semantics. Similarly, RSGPT [35] utilizes a Vicuna-v1.5-7B LLM, showing that while large language models generally yield improved performance in remote sensing tasks, considerations regarding training time and resource requirements remain central to practical deployments.
@@
Vision–Language Alignment: In the architecture of conversational Vision–Language Models, the middle connection layer plays a crucial role. It acts as a bridge between remote sensing images and large language models, projecting image information into a space that LLMs can comprehend, enabling them to understand complex remote sensing tasks. A common approach to bridge the modal gap is to use Multi-Layer Perceptrons (MLPs), which employ a single linear layer [54, 55] or double layers [56, 68, 99] to project visual tokens and align feature dimensions with word embeddings. Additionally, some works focus on designing connectors with intricate structures, offering valuable insights for future research.
SkyEyeGPT [54] is introduced as a unified multimodal large language model specifically designed for remote sensing vision–language understanding, with the objective of exploring the application of VLMs in the remote sensing field. The authors carefully create a remote sensing multimodal instruction-following dataset, which includes both single-task and multi-task dialogue instructions. After manual verification, the dataset consists of a high-quality remote sensing instruction-following corpus with a total of 968 K samples. SkyEyeGPT projects remote sensing visual features into the language domain via an alignment layer, and these features, along with task-specific instructions, are fed into a remote sensing decoder based on LLM to predict answers for open-ended remote sensing tasks. SkyEyeGPT employs a two-stage fine-tuning approach to enhance instruction-following and multi-turn dialogue capabilities at different granularities. Notably, it achieves excellent performance across a variety of tasks without requiring additional encoding modules.
@@
One reason current general VLMs perform poorly in the remote sensing field is that the visual encoder of LLMs usually uses CLIP, but CLIP’s ViT lacks spatial awareness, affecting VLM’s spatial understanding ability [102]. To address this issue, H2RSVLM [56] and SkySenseGPT [68] enhance the visual encoder’s perceptual ability at the source by creating higher-quality, large-scale, and detailed image–text pairs in remote sensing datasets. EarthGPT [55] expands the input images to include SAR and infrared images. Furthermore, some research endeavors opt for designing elaborate vision–language connection structures to enhance the multimodal perception capabilities of VLM. GeoChat [30] constructs bounding boxes for text representations to express their spatial coordinates, enabling it to perform object grounding or accept region inputs, thus having reasoning capabilities for remote sensing images. RSGPT [35], under the guidance of InstructBLIP [103], inserts an annotation-aware query transformer (Q-Former) to enhance the alignment of visual and textual features. Additionally, the annotation-aware mechanism interacts with query embeddings through self-attention layers, allowing RSGPT to extract comprehensive visual features crucial for RS vision–language tasks. LHRS-Bot [64] introduces a set of learnable queries for each level of image features. These queries aim to summarize the semantic information of each level of visual features through a series of stacked cross-attention and MLP layers.
@@
Model Year Enhancement Technique
@@
Txt2Img [104] 2023 Employs modern Hopfield layers for hierarchical prototype learning of text and image embeddings.
@@
CPSeg [105] 2024 Utilizes a “chain of thought” process that leverages textual information associated with the image.
@@
SHRNet [106] 2023 Utilizes a hash-based architecture to extract multi-scale image features.
@@
MGeo [107] 2023 Introduces the geographic environment (GC) extracting multimodal correlations for accurate query–POI matching.
@@
GeoCLIP [108] 2023 Via an image-to-GPS retrieval approach by explicitly aligning image features with corresponding GPS locations.
@@
SpectralGPT [109] 2024 Provides a foundational model for the spectral data based on the MAE [110] architecture.
@@
TEMO [111] 2023 Captures long-term dependencies from description sentences of varying lengths, generating text-modal features for each category.
5.3 Others Models
In addition to the aforementioned approaches, as shown in Table 4, several other advancements have been made to address domain-specific tasks in remote sensing.
Txt2Img [104] employs modern Hopfield layers for hierarchical prototype learning of text and image embeddings, enabling the generated remote sensing images to better preserve the shape information of ground objects while providing more detailed texture and boundary information. CPSeg [105] employs a novel “Chain-of-Thought” approach using language prompting to enhance semantic segmentation in remote sensing imagery, achieving superior performance, particularly in flood scenarios, through improved context-aware understanding and integration of textual descriptions. SHRNet [106] utilizes a hash-based architecture to extract multi-scale image features, while QRN improves alignment efficiency and enhances the visual–spatial reasoning ability of remote sensing Visual Question Answering systems. MGeo [107], a multimodal geographic language model, introduces the geographic environment (GC) as a new modality to enhance query–POI (Point of Interest) matching, outperforming baselines even without user geolocation, while its GC modeling holds potential for further improvements with extensions like POI images. GeoCLIP [108], the first CLIP-inspired image-to-GPS retrieval approach, introduces a novel Location Encoder that models the Earth as a continuous function using positional encoding and hierarchical representations. By aligning image features with GPS locations, GeoCLIP achieves competitive geo-localization performance, even with limited training data, and enables text-based geo-localization via its CLIP backbone. SpectralGPT [109], a spectral RS foundational model based on MAE [110], leverages 3D token generation and multi-target reconstruction to capture spatial–spectral patterns. Trained on one million spectral images with 600 M+ parameters, it achieves 99.21% for single-label Scene Classification on EuroSAT [112] and also demonstrates strong performance in semantic segmentation and Change Detection, advancing spectral RS big data applications. TEMO [111] introduces a text-modal knowledge extractor and a crossmodal assembly module to fuse text and visual features, reducing classification confusion in novel classes and enhancing few-shot object detection in remote sensing imagery. By leveraging a mask strategy, separation loss, and improved word embeddings to capture long-term dependencies, TEMO achieves 42.8%, 75.1% and 40.8% across DIOR [73], NWPU [113], and FAIR1M [72] in 10-shot scenarios, respectively.
@@ 5.4 Performance Comparison
In this section, we conduct a comparative analysis of recent Visual–Language Models, evaluating their performance on five widely recognized datasets and common visual–language tasks including Visual Question Answering and Image Captioning. VLMs have gained significant attention in the remote sensing domain over the past two years, with researchers exploring their applications in various tasks. However, models tailored for specialized domains, such as SpectralGPT [109] which is designed for spectral remote sensing, face challenges in cross-model comparisons due to unique datasets and tasks. Despite efforts to build versatile foundational models capable of handling diverse downstream tasks, the remote sensing field still lacks a unified benchmark for comprehensive performance evaluation. To address this gap, we focus on shared datasets and tasks to provide a consistent and meaningful comparison of VLM performance.
Table 5 summarizes the evaluation results for five prevalent datasets (e.g., AID [69], RSVGD [52]), highlighting the best-performing VLMs. Due to differing implementations of VLM pre-training methods, it is essential to identify standout models. Among the contrastive approaches, RemoteCLIP [29] emerges as a top performer, particularly excelling in WHU-RS19 [114]. Notably, GeoRSCLIP [31] demonstrates strong few-shot capabilities, achieving competitive results on NWPU-RESISC45 [70]. In comparison, conversational models, such as SkySenseGPT [68] and SkyEyeGPT [54], consistently achieve superior performance in multiple datasets, with SkySenseGPT leading in AID [69] and WHU-RS19 [114]. Unlike contrastive models, conversational methods also support Visual Grounding tasks, highlighting their broader applicability.
Table 5: Performance of VLM pre-training methods on common remote sensing datasets; bold indicates the best performance. $\ast$: few-shot prediction.
@@
Table 6: Performance of VLM pre-training methods on the Visual Question Answering task; bold indicates the best performance. $\ast$: few-shot prediction.
@@
SHRNet [106] 1 85.85 85.39 84.46
MQVQA [47] 1 - - 82.18
1 Only these are from the other part; the rest are conversational.
In Tables 6 and 7, we delve deeper into two frequently encountered visual–language tasks: Visual Question Answering (VQA) and Image Classification (IC). The results clearly illustrate that the conversational model, SkySenseGPT [68], maintains exceptional performance across both tasks. Notably, none of the contrastive VLMs proved effective for the VQA task, further underscoring the advantages of conversational methods. Additionally, in the IC task, the prevalence of conversational VLMs exceeds that of their contrastive counterparts.
In conclusion, from Tables 2 and 3, it is evident that a key architectural difference between conversational VLMs and contrastive VLMs lies in the use of pre-trained LLMs (e.g., Vicuna [78], LLaMA [20]) in conversational VLMs. These LLMs process visual elements extracted by visual encoders and transformed by middleware, along with user queries or textual instructions. This enables conversational VLMs to handle more diverse and complex visual–language tasks, such as VQA and IC. As shown in Tables 6 and 7, conversational VLMs consistently outperform other types, further unlocking the potential of language to enhance reasoning in complex scenarios. Consequently, we advocate for adopting conversational model architectures as the preferred choice for future advancements in visual–language modeling.
Table 7: Performance of VLM pre-training methods on image caption task; bold indicates the best performance.
@@
Syndney- Captions [85] CRSR 1 [92] 79.94 74.4 69.87 66.02 41.50 74.88 289.00
@@
1 Only CRSR is contrastive; the rest are conversational.
@@ 6 Conclusions and Future Work
This review provides a systematic overview of Vision–Language Models (VLM) in the field of remote sensing. It introduces milestone works in the development of VLM itself, discusses the current research status of VLM in remote sensing, outlines the tasks involved, and offers a brief comparison of different VLM methods in various remote sensing applications. Overall, VLM has once again facilitated advancements in data processing technologies within the field of remote sensing. Current work has clearly demonstrated the potential of VLM applications in this area. However, there are still some shortcomings in current VLM technologies and their applications in remote sensing. Future work could consider the following directions:
  • Addressing regression problems. Currently, the mainstream approach in VLM typically involves using a tokenizer to vectorize text, aligning the visual information with word vectors, and then feeding these into LLMs. However, this method fails to capture numerical relationships effectively, which leads to problems in regression tasks. For instance, a number like ‘100’ may be tokenized as ‘1’, ‘00’, or ‘100’, causing loss of precision and affecting regression accuracy. Future work should focus on enhancing VLMs for regression tasks by addressing this limitation. One promising direction is to develop specialized tokenizers for numerical values, which would ensure more accurate representation of numbers in text. Additionally, exploring the integration of task-specific regression heads into the existing VLM framework can help the model achieve a more efficient inference process. For example, REO-VLM [116] designs a regression head with a four-layer MLP-mixer-like structure that integrates knowledge embeddings derived from LLM hidden tokens with complex visual tokens. This innovative approach enables VLMs to effectively handle regression tasks such as Above-Ground Biomass (AGB) [117, 118] estimation in Earth observation (EO) applications. Task-specific regression heads enable the model to directly output data in formats that meet the specific requirements of the task, rather than being limited to regression on text tokens. Moreover, these heads can be dynamically adjusted based on the needs of different tasks, making the regression output more accurate and efficient. Introducing the Mixture of Experts (MOE) architecture to improve the adaptability and accuracy of regression tasks is also an excellent choice. When deployed in VLMs, it can use a gating mechanism to dynamically match the appropriate expert modules for different tasks, enabling the model to effectively handle multi-task learning with sufficient capability.
  • Aligning with the structural characteristics of remote sensing images. Current VLMs in remote sensing largely adopt frameworks developed for conventional computer vision tasks and rely on models pre-trained exclusively on RGB imagery, thereby failing to fully account for the distinctive structural and spectral characteristics of remote sensing data. This limitation is particularly evident for modalities such as SAR and HSI, where crucial features remain poorly captured by general-purpose extractors. As a result, while existing VLMs may perform reasonably well on RGB data, they struggle with SAR and HSI inputs and often depend on superficial fine-tuning with limited datasets. To address this gap, future research should emphasize both the design of specialized feature extractors tailored to the complexities of multispectral and SAR data, and the development of algorithms capable of handling multidimensional inputs. REO-VLM [116] makes effective attempts by introducing spectral recombination and pseudo-RGB strategies. It splits the high-dimensional MS image into multiple three-channel images and generates a third channel for the SAR images, which typically contain two polarization channels, thus creating a pseudo-RGB representation compatible with the pre-trained visual encoder. Expanding and diversifying remote sensing datasets to include large-scale image–text pairs from multiple modalities (e.g., RGB, SAR, HSI, LiDAR) will be essential for training VLMs that can achieve robust visual–text alignment across all data types. By integrating richer and more comprehensive feature representations, as well as exploring data fusion strategies, it will be possible to enhance VLM performance in a wide range of complex remote sensing applications, including land-use classification, urban monitoring, and disaster response.
  • Multimodal output. Early methods, such as classification models, were constrained by their output format and thus limited to a single task. Although most current VLMs, which produce text outputs, have broadened their scope and can effectively handle numerous tasks, they still fall short for dense prediction tasks like segmentation and Change Detection. To overcome this limitation, future research should focus on enabling VLMs to produce truly multimodal outputs—such as images, videos, and even 3D data. Recent studies have begun exploring the enhancement of traditional proprietary models by adding a text branch as an auxiliary component, thereby fostering richer multimodal feature interaction and improving image feature extraction and interpretation while preserving existing output heads. Building on this idea, another promising direction is to use a VLM as an agent that interfaces with multiple specialized model output heads and activates the most suitable one based on the task identified, thus effectively combining the reasoning power of LLMs with the expertise of dedicated models. In addition, a recent insightful approach is to use next-token prediction to achieve multimodal output. For example, Emu3 [119] adopts a completely next-token prediction-based architecture, avoiding the complex diffusion models or combinatorial methods used in the past. It innovatively uses visual tokens as part of the output to generate images and videos, and outperforms well-known open-source models.
  • Multitemporal Regression and Trend Inference. Current VLM research in the remote sensing field often focuses solely on static images, overlooking the potential of multitemporal data. Since remote sensing data can capture information that changes over time, future research should explore how to extend VLMs to handle multitemporal regression tasks. This involves not only spatial feature recognition but also the ability to infer temporal trends and dynamics. A key research direction is to develop methods that utilize historical remote sensing data to analyze and predict long-term environmental or ecological changes. This is particularly important for applications such as climate change monitoring, land-use change prediction, and ecosystem management, where temporal patterns play a crucial role. Remote sensing images with temporal continuity can be encoded as sequences, providing multitemporal data as input for training VLMs. By adding multitemporal regression capabilities to VLMs, the model could provide more context-aware predictions, improving the accuracy of long-term environmental assessments.